What is the magnitude of the columbic force between two point charges +3 microcoulumb and +5 microcoulumb separated from each other by a distance of 30cm in free space. What is the nature of the force?


1
Expert's answer
2021-05-24T15:44:22-0400

Gives

"q_1=3\mu c"

"q_2=5\mu c"

"r=30cm =0.3m"

"F=\frac{kq_1q_2}{r^2}"

"F=\frac{9\times10^9\times3\times10^{-6}\times5\times10^{-6}}{0.3^2}=1.5N"

Then

"q_1q_2>0"

Force is repulsive nature